# Learning the structure of any Hamiltonian from minimal assumptions

### v1

@@ 7       $\widehat{S}\leftarrow\widehat{S}\cup\{P^{(j)}\}$;
Let $\operatorname{\mathcal{A}}:(H,\epsilon,\delta)\mapsto\widehat{\lambda}\in\mathbb{R}^{m}$ be some base learning algorithm with the guarantee that $\Pr(\|\lambda-\widehat{\lambda}\|_{\infty}\leq\epsilon)\geq 1-\delta$. The idea is to bootstrap applications of this base algorithm, with constant precision $\epsilon_{0}=1/2$, to the Heisenberg limit $t_{\mathrm{total}}\sim 1/\epsilon$ by adaptively adjusting the Hamiltonian being learned via quantum control. Consider a previously learned estimate $\widehat{H}$ and form a “residual Hamiltonian” $R=H-\widehat{H}$. Assuming that the current estimate is only $\eta$-good, we can amplify $R$ by a factor of $1/\eta$ to boost the residual errors up to a constant. Then if we learn $R/\eta$ using $\operatorname{\mathcal{A}}$ with merely constant precision $\epsilon_{0}=1/2$, we can improve our estimate to $\eta/2$ error. Recursing on this process for multiple rounds $j=0,1,\ldots,T$ for some $T$, it is clear that each round produces an estimate with error $\eta_{j}=2^{-j}$. Hence $T=\lfloor\log_{2}(1/\epsilon)\rfloor$ suffices to get the target accuracy $\epsilon$. Taking some care with the acceptable amount of failure probability establishes the correctness of the algorithm. This overall strategy is outlined in Algorithm 2.
@@ 1.4 Discussion
    Is there an information-theoretic separation between Hamiltonian learning protocols with and without time reversal? While exponential separations have been established in unitary learning problems [cotler2023information, schuster2024random], these results do not immediately translate to our sparse Hamiltonian learning problem. Since we have established that this problem can be solved in polynomial time, any separation can only be polynomially large. Investigating the (im)possibility of nearly Heisenberg-limited scaling without time reversal would be an important question to address in this context.
@@
    Do nontrivial $m$-dependent lower bounds exist for 1.2? To our knowledge, the current best lower bound is $\Omega(1/\epsilon)$ [huang2023learning]. This bound is demonstrated by reducing the $m$-term Hamiltonian problem to distinguishing between two single-term Hamiltonians. Meanwhile, exponential lower bounds have been shown for protocols with ($\Omega(4^{n}/n)$) and without ($\Omega(4^{n}/\epsilon)$) ancillas and adaptivity, where now the distinguishing task corresponds to a net of Hamiltonians generated by Haar-random unitary transformations [bluhm2024hamiltonian]. Related is the problem of learning unitary channels, which has the tight lower bound of $\Omega(4^{n}/\epsilon)$ queries [haah2023query]. These results can be understood as either the case of constant $m$ or exponentially large $m$; a lower bound for the intermediate regime $m=\Theta(\operatorname{poly}(n))$ is of clear interest. Note added: During the final preparation stage of this manuscript, a lower bound of $\Omega(m/\epsilon_{1})$ with respect to $\ell_{1}$-error $\epsilon_{1}$ was claimed by Ma, Flammia, Preskill, and Tong [ma2024learning].
@@ 2.1 Notation
We denote the imaginary unit by $\mathrm{i}\equiv\sqrt{-1}$. For any integer $k\geq 1$, define the set $[k]\coloneqq\{1,\ldots,k\}$. The natural logarithm is denoted by $\log$, and other bases will be specified when necessary. Asymptotic upper and lower bounds are denoted by $\mathcal{O}(\cdot)$ and $\Omega(\cdot)$, respectively. We write $f=\Theta(g)$ if $f=\mathcal{O}(g)$ and $f=\Omega(g)$. Tildes on big-O notation suppress polylogarithmic factors, e.g., $\widetilde{\mathcal{O}}(f)=\mathcal{O}(f\operatorname{polylog}(f))$. When we only care that a quantity is polynomially bounded, we write $\operatorname{poly}(f,g,\ldots)=(fg\cdots)^{\mathcal{O}(1)}$. Generally speaking, for an object $z$, we will use $\widehat{z}$ to denote an estimate of it and $\widetilde{z}$ to denote an approximation to it.
@@ 2.2 Linear algebra
Let $\|x\|_{p}\coloneqq\mathopen{}\left(\sum_{j=1}^{d}|x_{j}|^{p}\right)^{1/p}\mathclose{}$ be the $\ell_{p}$-norm of any $x\in\mathbb{C}^{d}$. For all $0<r<p\leq\infty$, these vector norms obey
$\|x\|_{p}\leq\|x\|_{r}\leq d^{\frac{1}{r}-\frac{1}{p}}\|x\|_{p},$ (9)
with the understanding that $\frac{1}{\infty}=0$.
@@
When it is unimportant to specify the precise control bits $b$, or the context is otherwise clear, we will simply write $\texttt{ctrl}(U)$.
Finally, we will also work with quantum channels: completely positive, trace-preserving maps on the space of operators. The natural norm to equip quantum channels with is the diamond norm, also known as the completely bounded trace norm.
@@ 2.3 Block encodings
For brevity, we will say that $B$ is an $(\alpha,a,\varepsilon)$-BE. One may write an $(\alpha,a,0)$-BE as
@@
Let $y\in\mathbb{C}^{K}$ with $\|y\|_{1}\leq\beta$, and $\varepsilon\geq 0$. The pair of $b$-qubit unitaries $(\mathsf{PREP}_{L},\mathsf{PREP}_{R})$ is called a $(\beta,b,\varepsilon)$-state-preparation pair if
@@
Then $(\mathsf{PREP}_{L}^{\dagger}\otimes\mathbb{I}_{a+n})\mathsf{SEL}(\mathsf{PREP}_{R}\otimes\mathbb{I}_{a+n})$ is an $(\alpha\beta,a+b,\alpha\varepsilon_{1}+\beta\varepsilon_{2})$-block encoding of $A$ [^5] Note that a typo in [gilyen2019quantum] mistakenly carried an factor of $\alpha$ in the $\beta\varepsilon_{2}$ error term.
In this paper, we assume that we can produce state-preparation pairs with infinite precision. While some finite-precision machine error will be incurred in practice, this is vastly negligible to all other errors in our algorithm. Furthermore by the Solovay–Kitaev theorem, we only pay a polylogarithmic overhead for synthesizing the state-preparation pair (as well as any other quantum gates that we may need) from a finite gate set. If necessary, we trust the diligent reader to carry forth the machine precision throughout our exposition.
@@ 3.1 State preparation from time-evolution queries
[castaneda2023hamiltonian] defines the pseudo-Choi state with a reference control qubit as follows.
@@
[…]
@@
and $\max_{-\frac{1}{2}\leq x\leq\frac{1}{2}}|P(x)|\leq 1$. Thus, quantum signal processing or the quantum singular value transformation (on one ancilla qubit) can be applied to encode $P(\sin(H/\Delta))$. A polynomial transformation of degree $d$ requires $d$ queries to $U$ and $U^{\dagger}$ and $\mathcal{O}(d)$ additional quantum gates. This produces a $(\frac{\pi}{2},2,\frac{\pi}{2}\varepsilon^{\prime})$-BE of $H/\Delta$; letting $\varepsilon=\frac{\pi}{2}\varepsilon^{\prime}$ furnishes the claim. ∎
@@
This procedure uses $\mathcal{O}(\log(1/\varepsilon))$ queries to $U=e^{-\mathrm{i}H/\Delta}$, its inverse, and their multi-controlled forms.
@@ 3.2 Estimating with decoding operators
These operators are defined such that their mean values with respect to $|\Psi_{\mathrm{ref}}(\frac{\widetilde{H}}{\Delta\pi/2})\rangle$ recover $\widetilde{\lambda}_{a}$, up to normalizations:
@@
While $\Delta$ is chosen by the learner, $\mathcal{N}(\frac{\widetilde{H}}{\Delta\pi/2})$ is unknown. To access this quantity, $O_{\mathcal{N}}$ is used:
@@
Note that if we had not prepared the pseudo-Choi state with the reference control qubit, this normalization would have been more difficult to infer.
@@ 4 Identifying Pauli spectra with pseudo-Choi states
The Bell sampling algorithm can also discover elements $E_{a^{\prime}}\notin S_{\gamma}(\widetilde{H})$. At this current stage in the algorithm, we do not have enough information to determine inclusion or exclusion, so we store all elements observed. As long as we have at least a superset of $S_{\gamma}(\widetilde{H})$, we say that the structure learning algorithm has succeeded. This is fine because in the next stage, shadow tomography can easily estimate an overcomplete set of terms with only logarithmic cost to the sample complexity, per Eq. 34. Any terms that are too small can then be discarded. And because we only ever make a polynomial number of queries, we can never pick up more than a polynomial number of superfluous terms.
@@
Comparing this result to Proposition 3.5, we find remarkably that the query and evolution-time complexity of both parameter learning and structure learning from pseudo-Choi states are identical. Therefore even without invoking the precision bootstrap, we have demonstrated an algorithm that can perform structure learning on a completely arbitrary $n$-qubit Hamiltonian, using $\widetilde{\mathcal{O}}(\|H\|/\epsilon^{2})$ total evolution time.
To conclude this section, we make a remark regarding a potential constant shift in $H$.
@@ 5.3 Structure learning
To perform structure learning with (nearly) Heisenberg-limited scaling, the algorithm of Section 4 must also be placed into the bootstrap framework. We describe how that procedure works in this subsection.
@@ 6.1 Block encoding the matrix logarithm
Ideally, we would prefer to have constructed the LCU for $\log U$ using block encodings of $(U-\mathbb{I}_{n})^{k}$, wherein the coefficients satisfy $\sum_{k=1}^{K}|(-1)^{k+1}/k|=\mathcal{O}(\log K)$. However, each such block encoding is subnormalized by a factor of $2^{k}$, which again would require inverses to be boosted via singular value amplification [gilyen2019quantum].
@@ 6.3 Parameter estimation
This learns $\lambda_{a}^{\prime}$ to error $\epsilon$. Since $\lambda_{a}^{\prime}$ is $\gamma\Delta$-close to $\lambda_{a}$, as usual we take $\gamma=c\epsilon/\Delta$ for sufficiently small $c>0$ and improve our sampling error to $(1-c)\epsilon$ as well. This gets us a query complexity
@@ Acknowledgments
AZ thanks Andrew Baczewski, Matthias Caro, John Kallaugher, Danial Motlagh, and Ojas Parekh for helpful discussions. This work was supported by the Laboratory Directed Research and Development program at Sandia National Laboratories, under the Gil Herrera Fellowship in Quantum Information Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA-0003525. AZ also acknowledges support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Accelerated Research in Quantum Computing.

### v2

@@ 7       $\widehat{S}\leftarrow\widehat{S}\cup\{P^{(j)}\}$;
Let $\operatorname{\mathcal{A}}:(H,\epsilon,\delta)\mapsto\widehat{\lambda}\in\mathbb{R}^{m}$ be some base learning algorithm with the guarantee that $\Pr(\|\lambda-\widehat{\lambda}\|_{\infty}\leq\epsilon)\geq 1-\delta$. The idea is to bootstrap applications of this base algorithm, with constant precision $\epsilon_{0}=1/2$, to the Heisenberg limit $t_{\mathrm{total}}\sim 1/\epsilon$ by adaptively adjusting the Hamiltonian being learned via quantum control. Consider a previously learned estimate $\widehat{H}$ and form a “residual Hamiltonian” $R=H-\widehat{H}$. Assuming that the current estimate has error $\leq\eta$, we can amplify $R$ by a factor of $1/\eta$ to boost those residual errors up to a constant. Then if we learn $R/\eta$ using $\operatorname{\mathcal{A}}$ with merely constant precision $\epsilon_{0}=1/2$, we can improve our estimate to $\eta/2$ error. Recursing on this process for multiple rounds $j=0,1,\ldots,T$ for some $T$, it is clear that each round produces an estimate with error $\eta_{j}=2^{-j}$. Hence $T=\lfloor\log_{2}(1/\epsilon)\rfloor$ suffices to get the target accuracy $\epsilon$. Taking some care with the acceptable amount of failure probability establishes the correctness of the algorithm. This overall strategy is outlined in Algorithm 2.
@@ 1.4 Discussion
    Is there an information-theoretic separation between Hamiltonian learning protocols with and without time reversal? While exponential separations have been established in unitary learning problems [cotler2023information, schuster2024random], these results do not immediately translate to our sparse Hamiltonian learning problem. Since we have established that this problem can be solved in polynomial time, any separation can only be polynomially large.
@@
    What is the asymptotically tightest lower bound for the evolution time needed to solve 1.2? To our knowledge, the current best lower bound for $\ell_{\infty}$-error $\epsilon$ is $\Omega(1/\epsilon)$, obtained by Huang, Tong, Fang, and Su [huang2023learning] by reducing the $m$-term problem to a distinguishing task between two single-term Hamiltonians. For $\ell_{1}$-error $\epsilon_{1}$, Ma, Flammia, Preskill, and Tong [ma2024learning] proved that $\Omega(m/\epsilon_{1})$ evolution time is necessary. Meanwhile for non-sparse Hamiltonians ($m\leq 4^{n}$), Bluhm, Caro, and Oufkir [bluhm2024hamiltonian] showed exponential lower bounds for protocols with ($\Omega(4^{n}/n)$) and without ($\Omega(4^{n}/\epsilon)$) ancillas and adaptivity. A related problem is learning arbitrary unitary channels, whose query lower bound of $\Omega(4^{n}/\epsilon)$ was given by Haah, Kothari, O’Donnell, and Tang [haah2023query].
@@ 2.1 Notation
We denote the imaginary unit by $\mathrm{i}\equiv\sqrt{-1}$. For any integer $k\geq 1$, define the set $[k]\coloneqq\{1,\ldots,k\}$. The natural logarithm is denoted by $\log$, and other bases will be specified when necessary. Asymptotic upper and lower bounds are denoted by $\mathcal{O}(\cdot)$ and $\Omega(\cdot)$, respectively. We write $f=\Theta(g)$ if $f=\mathcal{O}(g)$ and $f=\Omega(g)$. Tildes on big-O notation suppress polylogarithmic factors, e.g., $\widetilde{\mathcal{O}}(f)\coloneqq\mathcal{O}(f\operatorname{polylog}(f))$. When we only care that a quantity is polynomially bounded, we write $\operatorname{poly}(f,g,\ldots)\coloneqq(fg\cdots)^{\mathcal{O}(1)}$. Generally speaking, for an object $z$, we will use $\widehat{z}$ to denote an estimate of it and $\widetilde{z}$ to denote an approximation of it.
@@ 2.2 Linear algebra
Let $\|x\|_{p}\coloneqq\mathopen{}\left(\sum_{j=1}^{d}|x_{j}|^{p}\right)^{1/p}\mathclose{}$ be the $\ell_{p}$-norm of any $x\in\mathbb{C}^{d}$, with $\|x\|_{\infty}\coloneqq\lim_{p\to\infty}\|x\|_{p}=\max_{j\in[d]}|x_{j}|$. For all $0<r<p$, these vector norms obey
$\|x\|_{p}\leq\|x\|_{r}\leq d^{\frac{1}{r}-\frac{1}{p}}\|x\|_{p}.$ (9)
These bounds also hold in the limit $p\to\infty$, in the sense that $\|x\|_{\infty}\leq\|x\|_{r}\leq d^{\frac{1}{r}}\|x\|_{\infty}$.
@@
When it is unimportant to specify the control bits $b$, or it is otherwise clear from context, we will simply write $\texttt{ctrl}(U)$.
Finally, we also work with quantum channels: completely positive, trace-preserving maps on the space of operators. The natural norm to equip quantum channels with is the diamond norm, also known as the completely bounded trace norm.
@@ 2.3 Block encodings
For brevity, we will use the shorthand: $(\alpha,a,\varepsilon)$-BE. One may write an $(\alpha,a,0)$-BE as
@@
Let $y\in\mathbb{C}^{K}$ with $\|y\|_{1}\leq\beta$, and $\varepsilon\geq 0$. The pair of $b$-qubit unitaries $(\mathsf{PREP}_{L},\mathsf{PREP}_{R})$ is called a $(\beta,b,\varepsilon)$-state-preparation pair for $y$ if
@@
Then $(\mathsf{PREP}_{L}^{\dagger}\otimes\mathbb{I}_{a+n})\mathsf{SEL}(\mathsf{PREP}_{R}\otimes\mathbb{I}_{a+n})$ is an $(\alpha\beta,a+b,\alpha\varepsilon_{1}+\beta\varepsilon_{2})$-block encoding of $A$ [^5] Note that a typo in [gilyen2019quantum] mistakenly carried a factor of $\alpha$ in the $\beta\varepsilon_{2}$ error term.
In this paper, we assume that we can produce state-preparation pairs with infinite precision. While some finite-precision machine error will be incurred in practice, this is vastly negligible to all other errors in our algorithm. Furthermore by the Solovay–Kitaev theorem, we only pay a polylogarithmic overhead for synthesizing the state-preparation pair (as well as any other quantum gates that we may need) from a finite gate set [dawson2006solovay]. If necessary, we trust the diligent reader to carry forth the machine precision throughout our exposition.
@@ 3.1 State preparation from time-evolution queries
Castaneda and Wiebe [castaneda2023hamiltonian] defined the pseudo-Choi state with a reference control qubit as follows.
@@
where the matrix representation on the rhs is in the |±$|\pm\rangle=\frac{1}{\sqrt{2}}(|0\rangle\pm|1\rangle)$ basis. This is a $(1,1,0)$-BE of $\sin(H/\Delta)$. Then, [gilyen2019quantum, Lemma 70] shows that there exists a polynomial $P(x)$ which approximates the $\arcsin(x)$ function on the domain $x\in[-\frac{1}{2},\frac{1}{2}]$. Specifically, for any $\varepsilon^{\prime}\in(0,\frac{1}{2}]$, this $P(x)$ is a polynomial of odd degree $d=\mathcal{O}(\log(1/\varepsilon^{\prime}))$ obeying
@@
and $\max_{-\frac{1}{2}\leq x\leq\frac{1}{2}}|P(x)|\leq 1$. Hence, we can apply the QSVT on a single ancilla qubit to encode $P(\sin(H/\Delta))$. A polynomial transformation of degree $d$ requires $d$ queries to $U$ and $U^{\dagger}$ and $\mathcal{O}(d)$ additional quantum gates. This produces a $(\frac{\pi}{2},2,\frac{\pi}{2}\varepsilon^{\prime})$-BE of $H/\Delta$; letting $\varepsilon=\frac{\pi}{2}\varepsilon^{\prime}$ furnishes the claim. ∎
@@
This procedure uses $\mathcal{O}(\log(1/\varepsilon))$ queries to $U=e^{-\mathrm{i}H/\Delta}$, its inverse, and their controlled forms.
@@ 3.2 Estimating with decoding operators
These operators are defined such that their mean values with respect to $|\Psi_{\mathrm{ref}}(\frac{\widetilde{H}}{\Delta\pi/2})\rangle$ recover $\widetilde{\lambda}_{a}$, up to normalizations: [^7] Note that the operators $O_{a}$ are not Hermitian, despite being constructed to have real-valued means. The postprocessing formulas used in classical shadows nonetheless easily accommodate complex-valued estimates.
@@
While $\Delta$ is chosen by the learner, $\mathcal{N}(\frac{\widetilde{H}}{\Delta\pi/2})$ is unknown. To access the value of this normalization factor, $O_{\mathcal{N}}$ is used:
@@
Note that if we had not prepared the pseudo-Choi state with the reference control qubit, this normalization factor would have been challenging to infer.
@@ 4 Identifying Pauli spectra with pseudo-Choi states
The Bell sampling algorithm can also discover elements $E_{a^{\prime}}\notin S_{\gamma}(\widetilde{H})$. At this current stage in the algorithm, we do not have enough information to determine inclusion or exclusion, so we store all elements observed. As long as we have at least a superset of $S_{\gamma}(\widetilde{H})$, we say that the structure learning algorithm has succeeded. In the next stage, shadow tomography can easily estimate an overcomplete set of terms with only a logarithmic sampling overhead, per Eq. 34. Any terms that are too small are then discarded; more precisely, if we set the shadow tomography error to $\epsilon/2$, then discarding all $|\widehat{\lambda}_{a}|\leq\epsilon$ ensures property 2 of 1.2. And because we only ever make a polynomial number of queries, we never identify more than a polynomial number of candidate terms.
@@
Comparing this result to Proposition 3.5, we find that the query and evolution-time complexity of both parameter learning and structure learning from pseudo-Choi states are identical. Therefore even without invoking the precision bootstrap, we have demonstrated an algorithm that can perform structure learning on a completely arbitrary $n$-qubit Hamiltonian, using $\widetilde{\mathcal{O}}(\|H\|/\epsilon^{2})$ total evolution time.
To conclude this section, we make a remark regarding the identity component in $H$.
@@ 5.3 Structure learning
To perform structure learning with Heisenberg-limited scaling (up to the $\log(1/\epsilon)$ factor), the algorithm of Section 4 must also be placed into the bootstrap framework. We describe how that procedure works in this subsection.
@@ 6.1 Block encoding the matrix logarithm
Ideally, we would prefer to have constructed the LCU for $\log U$ using block encodings of $(U-\mathbb{I}_{n})^{k}$, wherein the coefficients satisfy $\sum_{k=1}^{K}|(-1)^{k+1}/k|=\mathcal{O}(\log K)$. However, each such block encoding is subnormalized by a factor of $2^{k}$, which would then have to be boosted, say, using singular value amplification [gilyen2019quantum] which again requires $U^{\dagger}$.
@@ 6.3 Parameter estimation
This learns $\lambda_{a}^{\prime}$ to error $\epsilon$. Since $\lambda_{a}^{\prime}$ is $(\gamma\Delta)$-close to $\lambda_{a}$, as usual we take $\gamma=c\epsilon/\Delta$ for sufficiently small $c>0$ and improve our sampling error to $(1-c)\epsilon/2$ as well. This gets us a query complexity
@@ Acknowledgments
AZ thanks Andrew Baczewski, Matthias Caro, John Kallaugher, Danial Motlagh, and Ojas Parekh for helpful discussions, as well as multiple anonymous reviewers for their thorough feedback. This work was supported by the Laboratory Directed Research and Development program at Sandia National Laboratories, under the Gil Herrera Fellowship in Quantum Information Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA-0003525. AZ also acknowledges support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Accelerated Research in Quantum Computing.